Can you please help me with a problem? i have to identify the conic section and write the standard form and then graph it. x^2+y^2-4y-5=0 Thank you!
Answer by lwsshak3(11628)

Can you please help me with a problem? i have to identify the conic section and write the standard form and then graph it. x^2+y^2-4y-5=0
**
x^2+y^2-4y-5=0
complete the square
x^2+(y^2-4y+4)=5+4=9
x^2+(y-2)^2=9
This is an equation of a circle of standard form: (x-h)^2+(y-k)^2=r^2, (h,k)=center, r=radius.
For given equation:
center: (0,2)
radius=√9=3
see graph below:
..
y=±